40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 294     Entry time: Wed Aug 18 20:11:44 2010
Author: Frank 
Type: Summary 
Category: DAQ 
Subject: some signals moved to VME DAQ system 

moved the first 8 signals, the 4 individual sensors for each cavity, to the VME DAQ system.
Because i would have to open the foam box in order change the cables, i extended the existing ones for now to see if everything is working.
Once all channels are configured and everything is working i will open the box and make all the changes.
I also have to add the voltage reference for the second board and add the averaged channel from that board to the DAQ as well.
I need two more of the 9-pin D-SUB connector blocks or i have to make breakout cables instead.

Also added the VCO input monitor signal to the DAQ: channel name is C3:PSL-ACAV_VCOMON.
This channel can be used to keep the VCO signal centered in the range by feeding back to the temp setpoint of the cavity or the laser temp if locked alone (to match resonance of both cavities)

ELOG V3.1.3-